OpenStudy (anonymous):

Bob has two savings accounts. He deposited $200 more into account A than account B. After a period of time, account A has earned $84 in interest at 3%, and account B has earned $24 in interest at a rate of 2%. Write an equation to represent the situation. Explain each variable used. How much money did Bob initially deposit into each account? Solve the equation.

OpenStudy (anonymous):

I assume you mean Account A earned 3% annually and Account B earned 2% annually?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

OK. Let's say P is the amount Bob deposited into Account B. Then P+200 is the amount Bob deposited into Account A. Both accounts are allowed to sit for a period of t years, but Account A makes 3% interest, and Account B only makes 2% interest. So do you know the equation for interest compounded annually? It's \[A=P(1+r)^t\] whereas A is the amount you get in total, P is your principal (your initial deposit), r is your interest rate in decimal form and t is your time in years. So \[(P+200)(1+0.03)^t=P+200+84\] \[P(1+0.02)^t=P+24\] Now just solve this system and you will find P, the initial amount Bob deposited to account B.